in a rectangular garden whose area is 24 centimeters, the length is 5 centimeters more than the width. find the dimensions of the garden.
Answer by Nate(3500)

length = l = 5+w
width = w = w
area = a = 24
----------------------------------------------------------------------------
a=lw
24=w(w+5)
w^2+5w-24=0
(w+8)(w-3)=0
w=-8 or w=3
since you cannot have a negative length, your width is 3 and your length is 8!